A manufacturer of light bulbs produces bulbs that last a mean of 950 hrs. with a standard deviation of 120 hrs.What is the probability that the mean lifetime of random sample of 10 of these bulbs is less than 900 hours?
1
Expert's answer
2020-03-11T14:20:16-0400

Let "X" be the random variable which represents the length of the life of a light bulb, in hours: "X\sim N(\mu, \sigma^2/n)"

Then


"Z={X-\mu \over \sigma/\sqrt{n}}\sim N(0,1)"

Given that "\mu=950\ hrs, \sigma=120\ hrs, n=10."

The probability that the mean lifetime of random sample of 10 of these bulbs is less than 900 hours is


"P(X<900)=P(Z<{900-950 \over 120/\sqrt{10}})\approx P(Z<-1.3176)\approx"

"\approx0.0938"